OpenStudy (anonymous):

Suppose that m varies inversely as the cube of n. When n = 4, m = 0.25. Find m when n = 2.

hero (hero):

In general when y varies inversely as x, y = k/x or xy = k In this case, m varies inversely as the cube of n, therefore, m = k/n^3 or n^3m = k since n = 4 and m = 0.25, we substitute that into the equation to get 4^3 * .25 = k, then solve for k: 64*.25 = k 16 = k Do you believe you can find m when n = 2 now?